How many phone numbers are there with the code 270?
Answer by jim_thompson5910(35256)

Assuming you can use any digit in anyway you want, there are 10^7 = 10,000,000 (ten million) phone numbers.

If you're not allowed numbers like (270) 000-5555 or (270) 000-1234, then....

There are 9 numbers to choose from for the first digit and 10 numbers to choose from for the remaining slots. So there are 9*10^6 = 9,000,000 (nine million) phone numbers possible. This eliminates numbers like (270) 000-5555 or (270) 000-1234 but allows for numbers like (270) 300-5555 or (270) 300-1234

If you need to add more restrictions, then let me know.

Note: all of these phone numbers are completely fake and made up. No such real phone numbers exist (I don't think...)